a point charge develpos an electric field pf 40N/C and a potential diffrence of 10/C at a point .calculate the magnitude of charge and distance of point charge

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
Given information : 
 
Electric Field = 40 N/C 
 
Potential Difference = 10/C 
 
As we realize that 
 
Electric field = kq/r^2 
 
Potential distinction = kq/r 
 
Electric Field/potential distinction = kq/r^2/kq/r 
 
40/10 = 1/r 
 
r = ¼ = 0.25 m 
 
As we definitely realize that 
 
V = kq/r 
 
10 = 36 * 10^9 *q 
 
q = 10^-8/36 
 
q = 2 * 10^-10 C .
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)